# Origin of effective graviton mass from spontaneous symmetry breaking of scalar field

### v1

@@ Appendix C Linearised massive gravity theory
where $m$ is the graviton mass, $h_{\mu\nu}$ is rank-2, symmetric tensor field defined about a D-dimensional Minkowski space-time $\eta_{\mu\nu}=\text{diag}(-1,1,\ldots,1)$, $h=\eta^{\mu\nu}h_{\mu\nu}$ is the trace of the metric perturbation, $\alpha$ is referred to as the massive deformation parameter [ref] and $\mathscr{L}_{m=0}$ is the Lagrangian density corresponding to the linearized Einstein-Hilbert action, which is given by:
$\mathscr{L}_{m=0}=-\frac{1}{2}\partial_{\lambda}h_{\mu\nu}\partial^{\lambda}h^{\mu\nu}+\partial_{\mu}h_{\nu\lambda}\partial^{\nu}h^{\mu\lambda}-\partial_{\mu}h^{\mu\nu}\partial_{\nu}h+\frac{1}{2}\partial_{\lambda}h\partial^{\lambda}h\,.$ (118)
Setting $m=0$ in the above action leads to the linearized Einstein-Hilbert action, which describes a massless helicity-2 particle (massless graviton) with the gauge symmetry under the following transformation:
@@
$\xi^{\mu}$ is a spacetime-dependent infinitesimal gauge transformation. In the rest of this section, we will explore implications for two values of $\alpha$: $\alpha=1$ and $\alpha\neq 1$. Setting $\alpha=1$ in the above action leads to the Fierz-Pauli action [ref], while $\alpha\neq 1$ leads to dynamical $h$. We will also consider a special case $\alpha=1/2$, which we refer to as extended Fierz-Pauli action.
C.1 Fierz-Pauli action ($\alpha=1$)
@@
Although the above three equations are exactly equivalent to (120), it is easier to count the number of degrees of freedom (DoF) by using the above set of equations. The first equation is the evolution of a symmetric tensor field $h_{\mu\nu}$ and implies $10$ DoF in $D=4$. The traceless condition reduces one degree of freedom by forming a single constraint on the system. The transverse condition leads to $4$ more constraints. Hence, the Fierz-Pauli action consists of $5$ DoF in $D=4$ spacetime dimension.
C.2 Generalized Fierz-Pauli action ($\alpha\neq 1$)
@@
C.3 Extended Fierz-Pauli action ($\alpha=1/2$) in 4-D
@@
Appendix D Stueckelberg mechanism in linearised massive gravity theory
@@
D.1 Fierz-Pauli action ($\alpha=1$)
@@
At this point, one might think to consider scaling $V_{\mu}\rightarrow V_{\mu}/m$ to normalize the vector kinetic term, then take the $m\rightarrow 0$ limit. However, in this scenario, the resulting massless graviton and massless photon would yield 4 DoF (in 4 dimensions). Hence, $m\rightarrow 0$ is singular, leading to the loss of one of the original 5 DoF.
@@
$\begin{split}S&=\int d^{D}x\Big{[}\mathscr{L}_{m=0}-\frac{1}{2}F_{\mu\nu}F^{\mu\nu}-2(h_{\mu\nu}\partial^{\mu}\partial^{\nu}\phi-h\Box\phi)+\kappa h_{\mu\nu}T^{\mu\nu}\Big{]}\,.\end{split}$ (142)
@@
$\mathscr{L}_{m=0}[h]=\mathscr{L}_{m=0}[h^{\prime}]+(D-2)\Big{[}\partial_{\mu}\pi\partial^{\mu}h^{\prime}-\partial_{\mu}\pi\partial_{\nu}h^{{}^{\prime}\mu\nu}+\frac{1}{2}(D-1)\partial_{\mu}\pi\partial^{\mu}\pi\Big{]}.$ (144)
@@
$S=\int d^{D}x\Big{[}\mathscr{L}_{m=0}[h^{\prime}]-\frac{1}{2}F_{\mu\nu}F^{\mu\nu}-\frac{D-1}{D-2}\partial_{\mu}\phi\partial^{\mu}\phi+\kappa h_{\mu\nu}^{\prime}T^{\mu\nu}+\frac{2}{D-2}\kappa\phi T\Big{]},$ (145)
@@
It is easy to see that in $D=4$, the above action has five DoF —- two DoF carried by the canonical massless graviton, two DoF carried by the canonical massless vector, and one DoF carried by the canonical massless scalar. To confirm or infirm that these extra DoF do not introduce ghost degrees of freedom causing instabilities, we return to the action (138).
@@
$\begin{split}S&=\int d^{D}x\Big{[}\mathscr{L}_{m=0}[h^{\prime}]-\frac{1}{2}m^{2}(h_{\mu\nu}^{\prime}h^{{}^{\prime}\mu\nu}-h^{{}^{\prime}2})-\frac{1}{2}F_{\mu\nu}F^{\mu\nu}+2\frac{D-1}{D-2}\phi\left(\Box+\frac{D}{D-2}m^{2}\right)\phi\\ &-2m\left(h_{\mu\nu}^{\prime}\partial^{\mu}V^{\nu}-h^{\prime}\partial_{\mu}V^{\mu}-2\frac{D-1}{D-2}\phi\partial_{\mu}V^{\mu}\right)+2\frac{D-1}{D-2}m^{2}h^{\prime}\phi+\kappa h_{\mu\nu}^{\prime}T^{\mu\nu}+\frac{2}{D-2}\phi T\Big{]}.\end{split}$ (147)
@@
$\begin{split}S+S_{GF1}+S_{GF2}&=\int d^{D}x\Big{[}\frac{1}{2}h_{\mu\nu}^{\prime}(\Box-m^{2})h^{{}^{\prime}\mu\nu}-\frac{1}{4}h^{\prime}(\Box-m^{2})h^{\prime}+V_{\mu}(\Box-m^{2})V^{\mu}\\ &+2\frac{D-1}{D-2}\phi(\Box-m^{2})\phi+\kappa h_{\mu\nu}^{\prime}T^{\mu\nu}+\frac{2}{D-2}\kappa\phi T\Big{]}.\end{split}$ (150)
@@
$\begin{split}\frac{-i}{p^{2}+m^{2}}\Big{[}\frac{1}{2}(\eta_{\alpha\sigma}\eta_{\beta\lambda}+\eta_{\alpha\lambda}\eta_{\beta\sigma})-\frac{1}{D-2}\eta_{\alpha\beta}\eta_{\sigma\lambda}\Big{]},\ \frac{1}{2}\frac{-i\eta_{\mu\nu}}{p^{2}+m^{2}},\ \frac{(D-2)}{4(D-1)}\frac{-i}{p^{2}+m^{2}}.\end{split}$ (151)
@@
D.2 Extended Fierz-Pauli action ($\alpha=1/2$) in 4-D
@@
$\begin{split}-\frac{1}{2}m^{2}\left(h_{\mu\nu}h^{\mu\nu}-\frac{1}{2}h^{2}\right)&=-\frac{1}{2}m^{2}\Big{[}h_{\mu\nu}^{(T)}h^{(T)\mu\nu}+2h_{\mu\nu}^{(T)}(\partial^{\mu}h^{(L)\nu}+\partial^{\nu}h^{(L)\mu})+2\partial_{\mu}h_{\nu}^{(L)}\partial^{\mu}h^{(L)\nu}\\ &+2\partial_{\mu}h_{\nu}^{(L)}\partial^{\nu}h^{(L)\mu}-\frac{1}{2}(h^{(T)2}+4h^{(T)}\partial_{\mu}h^{(L)\mu}+4(\partial_{\mu}h^{(L)\mu})^{2})\Big{]}.\end{split}$ (153)
@@
$-\frac{1}{2}m^{2}(4\partial_{\mu}\partial_{\nu}l^{\parallel}\partial^{\mu}\partial^{\nu}l^{\parallel}-2(\Box l^{\parallel})^{2})=-m^{2}(\Box l^{\parallel})^{2}=-m^{2}l^{\parallel}(\Box)^{2}l^{\parallel}.$ (155)
@@
$\begin{split}-\frac{1}{2}m^{2}\left(h_{\mu\nu}h^{\mu\nu}-\frac{1}{2}h^{2}\right)&=-\frac{1}{2}m^{2}\Big{[}h_{\mu\nu}^{(T)}h^{(T)\mu\nu}-\frac{1}{2}h^{(T)2}-4h^{(L)\nu}\left(\partial^{\mu}h_{\mu\nu}^{(T)}-\frac{1}{2}\partial_{\nu}h^{(T)}\right)\\ &+2\partial_{\mu}l_{\nu}^{\perp}\partial^{\mu}l^{\perp\nu}-2m^{2}l^{\parallel}(\Box)^{2}l^{\parallel}\Big{]}\\ &=-\frac{1}{2}m^{2}\Big{[}h_{\mu\nu}^{(T)}h^{(T)\mu\nu}-\frac{1}{2}h^{(T)2}+2\partial_{\mu}l_{\nu}^{\perp}\partial^{\mu}l^{\perp\nu}+2l^{\parallel}(\Box)^{2}l^{\parallel}\Big{]},\end{split}$ (156)
@@
$\begin{split}S&=\int d^{4}x[\mathscr{L}_{FP}+\mathscr{L}_{1}],\ \mathscr{L}_{1}=-\frac{1}{4}m^{2}h^{2}\\ &=\int d^{4}x\Big{[}-\frac{1}{2}\partial_{\lambda}h_{\mu\nu}\partial^{\lambda}h^{\mu\nu}+\partial_{\mu}h_{\nu\lambda}\partial^{\nu}h^{\mu\lambda}-\partial_{\mu}h^{\mu\nu}\partial_{\nu}h+\frac{1}{2}\partial_{\lambda}h\partial^{\lambda}h\\ &-\frac{1}{2}m^{2}\left(h_{\mu\nu}h^{\mu\nu}-h^{2}\right)-\frac{1}{4}m^{2}h^{2}\Big{]}\,.\end{split}$ (157)
@@
$\mathscr{L}_{1}=-\frac{1}{4}m^{2}\Big{[}h^{2}+4h\partial_{\mu}V^{\mu}+4(\partial_{\mu}V^{\mu})^{2}+4h\Box\phi+8\partial_{\mu}V^{\mu}\Box\phi+4(\Box\phi)^{2}\Big{]}.$ (159)
@@
$\mathscr{L}_{1}=-\frac{1}{4}m^{2}h^{2}-mh\partial_{\mu}V^{\mu}-(\partial_{\mu}V^{\mu})^{2}-h\Box\phi-2\frac{1}{m}\partial_{\mu}V^{\mu}\Box\phi-\frac{1}{m^{2}}(\Box\phi)^{2}.$ (160)
@@
$\begin{split}\mathscr{L}_{1}&=-\frac{1}{4}m^{2}h^{{}^{\prime}2}-mh^{\prime}\partial_{\mu}V^{\mu}-(\partial_{\mu}V^{\mu})^{2}-h^{\prime}\Box\phi-2\frac{1}{m}\partial_{\mu}V^{\mu}\Box\phi-\frac{1}{m^{2}}(\Box\phi)^{2}\\ &-2m^{2}h^{\prime}\pi-4m^{2}\pi^{2}-4m\pi\partial_{\mu}V^{\mu}-4\pi\Box\phi.\end{split}$ (161)
@@
$\begin{split}\mathscr{L}_{1}&=-\frac{1}{4}m^{2}h^{{}^{\prime}2}-mh^{\prime}\partial_{\mu}V^{\mu}-(\partial_{\mu}V^{\mu})^{2}-h^{\prime}\Box\phi-2\frac{1}{m}\partial_{\mu}V^{\mu}\Box\phi-\frac{1}{m^{2}}(\Box\phi)^{2}\\ &-2m^{2}h^{\prime}\phi-4m^{2}\phi^{2}-4m\phi\partial_{\mu}V^{\mu}-4\phi\Box\phi.\end{split}$ (162)
@@
$\begin{split}S&=\int d^{4}x\Big{[}\mathscr{L}_{m=0}[h^{\prime}]-\frac{1}{2}m^{2}(h_{\mu\nu}^{\prime}h^{{}^{\prime}\mu\nu}-h^{{}^{\prime}2})-\frac{1}{2}F_{\mu\nu}F^{\mu\nu}+\left(-\phi\Box\phi+2m^{2}\phi^{2}\right)\\ &-2m(h_{\mu\nu}^{\prime}\partial^{\mu}V^{\nu}-h^{\prime}\partial_{\mu}V^{\mu})+(m^{2}h^{\prime}\phi+2m\phi\partial_{\mu}V^{\mu})+\kappa h_{\mu\nu}^{\prime}T^{\mu\nu}+\kappa\phi T\\ &-\frac{1}{4}m^{2}h^{{}^{\prime}2}-mh^{\prime}\partial_{\mu}V^{\mu}-(\partial_{\mu}V^{\mu})^{2}-h^{\prime}\Box\phi-2\frac{1}{m}\partial_{\mu}V^{\mu}\Box\phi-\frac{1}{m^{2}}(\Box\phi)^{2}\Big{]}\,.\end{split}$ (163)
@@
$\begin{split}S+S_{GF1}+S_{GF2}&=\int d^{4}x\Big{[}\mathscr{L}_{m=0}[h^{\prime}]-\frac{1}{2}m^{2}(h_{\mu\nu}^{\prime}h^{{}^{\prime}\mu\nu}-h^{{}^{\prime}2})-\frac{1}{2}F_{\mu\nu}F^{\mu\nu}+\left(-\phi\Box\phi+3m^{2}\phi^{2}\right)\\ &-\left(\partial^{\nu}h_{\mu\nu}^{\prime}-\frac{1}{2}\partial_{\mu}h^{\prime}\right)^{2}-m^{2}V_{\mu}V^{\mu}+\kappa h_{\mu\nu}^{\prime}T^{\mu\nu}+\phi T\Big{]}\\ &=\int d^{4}x\Big{[}\frac{1}{2}h_{\mu\nu}^{\prime}(\Box-m^{2})h^{{}^{\prime}\mu\nu}-\frac{1}{4}h^{\prime}(\Box-m^{2})h^{\prime}+V_{\mu}(\eta^{\mu\nu}\Box-\partial^{\mu}\partial^{\nu}-m^{2}\eta^{\mu\nu})V_{\nu}\\ &-3\phi(\Box-m^{2})\phi+\kappa h_{\mu\nu}^{\prime}T^{\mu\nu}+\kappa\phi T\Big{]}.\end{split}$ (165)
The relative sign of the first and third terms in the last line of the above equation does not change the equations of motion. Therefore, classically, such a change in sign is valid. Moreover, the above alterations of two signs do not change the theory even at the quantum level since all the correlation functions in quantum field theory that come from a single generating functional remain unchanged. The above action suggests no ghost DoF for $\alpha=1/2$ in four dimensions. However, for an arbitrary $\alpha$ in general $D$-dimesional space-time, the coefficients of each term may not be such that we can absorb the $(\Box\phi)^{2}$ in the gauge-fixing action.
Appendix E Degravitation
As mentioned in the main text, the graviton mass can be interpreted as a high-pass filter that sources must pass through before being detected by the graviton [66, 64, 108]. This is analogous to the propagation of electromagnetic waves in plasma. In unmagnetized neutral plasma, the dispersion relation for electromagnetic waves is $\omega^{2}=k^{2}+\omega_{p}^{2}$, where $\omega_{p}$ is plasma frequency. Thus, waves with frequencies lower than $\omega_{p}$ cannot propagate in plasma. In the case of massive gravity, the short wavelength has little effect. As the cosmological constant is inherently a constant, it operates like a source with an exceedingly long wavelength. Thus, the cosmological constant acts as a screen that serves as a high-pass filter, leading to a considerably diminished observed effective value. This process is commonly referred to as degravitation [66, 64, 108].

### v2

@@ Appendix B Fierz-Pauli action: Standard, Generalized and Extended
where $m$ is the mass of spin-2 field, $h_{\mu\nu}$ is rank-2, symmetric tensor field defined about a D-dimensional Minkowski space-time $\eta_{\mu\nu}=\text{diag}(-1,1,\ldots,1)$, $h=\eta^{\mu\nu}h_{\mu\nu}$ is the trace of the metric perturbation, $\alpha$ is referred to as the massive deformation parameter and $\mathscr{L}_{m=0}$ is the Lagrangian density corresponding to the linearized Einstein-Hilbert action, which is given by:
$\begin{split}\mathscr{L}_{m=0}&=-\frac{1}{2}\partial_{\lambda}h_{\mu\nu}\partial^{\lambda}h^{\mu\nu}+\partial_{\mu}h_{\nu\lambda}\partial^{\nu}h^{\mu\lambda}-\partial_{\mu}h^{\mu\nu}\partial_{\nu}h\\ &+\frac{1}{2}\partial_{\lambda}h\partial^{\lambda}h\,.\end{split}$ (68)
Setting $m=0$ in the above action leads to the linearized Einstein-Hilbert action, which describes a massless helicity-2 particle (massless spin-2 field) with the gauge symmetry under the following transformation:
@@
$\xi^{\mu}$ is a spacetime-dependent infinitesimal gauge transformation. In the rest of this section, we will explore implications for two values of $\alpha$: $\alpha=1$ and $\alpha\neq 1$. Setting $\alpha=1$ in the above action leads to the Fierz-Pauli action, while $\alpha\neq 1$ leads to dynamical $h$. We will also consider a special case $\alpha=1/2$, which we refer to as extended Fierz-Pauli action.
B.1 Fierz-Pauli action ($\alpha=1$)
@@
Although the above three equations are exactly equivalent to (70), it is easier to count the number of DoF by using the above set of equations. The first equation is the evolution of a symmetric tensor field $h_{\mu\nu}$ and implies $10$ DoF in $D=4$. The traceless condition reduces one degree of freedom by forming a single constraint on the system. The transverse condition leads to $4$ more constraints. Hence, the Fierz-Pauli action consists of $5$ DoF in $D=4$ spacetime dimension.
One of the striking features of the massive Spin-2 field, compared to massive spins $0,1/2,1,$ and $3/2$ fields, is that one can not straightforwardly extend the framework from Minkowski to arbitrary curved space-times [96, 97, 98, 99, 100]. As can be seen from Eq. (59), describing the massive spin-2 field in curved space-time depends on the background Ricci tensor and Ricci scalar. While the Fierz-Pauli action in Minkowski space-time has only five dynamical polarizations, it has six in arbitrary curved space-time. The additional polarization is a ghost [96, 97, 98, 99, 100]. As can be seen from Eq.  (59), the Fierz-Pauli theory for a free massive spin-2 field can be described in Einstein spaces, defined by the relation $R_{\mu\nu}=\Lambda g_{\mu\nu}$.
Generally, it is believed that $\alpha\neq 1$ in the massive gravity theory leads to ghost-like DoF with negative norm states. In our analysis of the effect of SSB on the dynamics of the spin-2 field, we find a massive gravity theory with extended Fierz-Pauli action where the massive deformation parameter takes the value $\alpha=1/2$. In the next two sections, we have shown explicitly that $\alpha=1/2$ also does not generate ghost DoF in generalized massive gravity theory.
B.2 Generalized Fierz-Pauli action ($\alpha\neq 1$)
@@
B.3 Extended Fierz-Pauli action ($\alpha=1/2$) in 4-D
@@
Appendix C Stueckelberg mechanism in linearised massive gravity theory
@@
C.1 Fierz-Pauli action ($\alpha=1$)
@@
At this point, one might think to consider scaling $V_{\mu}\rightarrow V_{\mu}/m$ to normalize the vector kinetic term, then take the $m\rightarrow 0$ limit. However, in this scenario, the resulting massless spin-2 field and massless photon would yield 4 DoF (in 4 dimensions). Hence, $m\rightarrow 0$ is singular, leading to the loss of one of the original 5 DoF.
@@
$\begin{split}S&=\int d^{D}x\Big{[}\mathscr{L}_{m=0}-\frac{1}{2}F_{\mu\nu}F^{\mu\nu}-2(h_{\mu\nu}\partial^{\mu}\partial^{\nu}\phi-h\Box\phi)\\ &+\kappa h_{\mu\nu}T^{\mu\nu}\Big{]}\,.\end{split}$ (92)
@@
$\begin{split}\mathscr{L}_{m=0}[h]&=\mathscr{L}_{m=0}[h^{\prime}]+(D-2)\Big{[}\partial_{\mu}\pi\partial^{\mu}h^{\prime}-\partial_{\mu}\pi\partial_{\nu}h^{{}^{\prime}\mu\nu}\\ &+\frac{1}{2}(D-1)\partial_{\mu}\pi\partial^{\mu}\pi\Big{]}.\end{split}$ (94)
@@
$\begin{split}S&=\int d^{D}x\Big{[}\mathscr{L}_{m=0}[h^{\prime}]-\frac{1}{2}F_{\mu\nu}F^{\mu\nu}-\frac{D-1}{D-2}\partial_{\mu}\phi\partial^{\mu}\phi\\ &+\kappa h_{\mu\nu}^{\prime}T^{\mu\nu}+\frac{2}{D-2}\kappa\phi T\Big{]},\end{split}$ (95)
@@
It is easy to see that in $D=4$, the above action has five DoF —- two DoF carried by the canonical massless spin-2 field, two DoF carried by the canonical massless vector, and one DoF carried by the canonical massless scalar. To confirm or infirm that these extra DoF do not introduce ghost DoF causing instabilities, we return to the action (88).
@@
$\begin{split}S&=\int d^{D}x\Big{[}\mathscr{L}_{m=0}[h^{\prime}]-\frac{1}{2}m^{2}(h_{\mu\nu}^{\prime}h^{{}^{\prime}\mu\nu}-h^{{}^{\prime}2})-\frac{1}{2}F_{\mu\nu}F^{\mu\nu}\\ &+2\frac{D-1}{D-2}\phi\left(\Box+\frac{D}{D-2}m^{2}\right)\phi\\ &-2m\left(h_{\mu\nu}^{\prime}\partial^{\mu}V^{\nu}-h^{\prime}\partial_{\mu}V^{\mu}-2\frac{D-1}{D-2}\phi\partial_{\mu}V^{\mu}\right)\\ &+2\frac{D-1}{D-2}m^{2}h^{\prime}\phi+\kappa h_{\mu\nu}^{\prime}T^{\mu\nu}+\frac{2}{D-2}\phi T\Big{]}.\end{split}$ (97)
@@
$\begin{split}S&+S_{GF1}+S_{GF2}=\int d^{D}x\Big{[}\frac{1}{2}h_{\mu\nu}^{\prime}(\Box-m^{2})h^{{}^{\prime}\mu\nu}-\frac{1}{4}h^{\prime}(\Box-m^{2})h^{\prime}\\ &+V_{\mu}(\Box-m^{2})V^{\mu}+2\frac{D-1}{D-2}\phi(\Box-m^{2})\phi+\kappa h_{\mu\nu}^{\prime}T^{\mu\nu}+\frac{2}{D-2}\kappa\phi T\Big{]}.\end{split}$ (100)
@@
$\begin{split}\frac{-i}{p^{2}+m^{2}}&\Big{[}\frac{1}{2}(\eta_{\alpha\sigma}\eta_{\beta\lambda}+\eta_{\alpha\lambda}\eta_{\beta\sigma})-\frac{1}{D-2}\eta_{\alpha\beta}\eta_{\sigma\lambda}\Big{]},\ \frac{1}{2}\frac{-i\eta_{\mu\nu}}{p^{2}+m^{2}},\\ &\frac{(D-2)}{4(D-1)}\frac{-i}{p^{2}+m^{2}}.\end{split}$ (101)
@@
C.2 Extended Fierz-Pauli action ($\alpha=1/2$) in 4-D
@@
$\begin{split}-\frac{1}{2}m^{2}&\left(h_{\mu\nu}h^{\mu\nu}-\frac{1}{2}h^{2}\right)=-\frac{1}{2}m^{2}\Big{[}h_{\mu\nu}^{(T)}h^{(T)\mu\nu}\\ &+2h_{\mu\nu}^{(T)}(\partial^{\mu}h^{(L)\nu}+\partial^{\nu}h^{(L)\mu})+2\partial_{\mu}h_{\nu}^{(L)}\partial^{\mu}h^{(L)\nu}\\ &+2\partial_{\mu}h_{\nu}^{(L)}\partial^{\nu}h^{(L)\mu}-\frac{1}{2}(h^{(T)2}+4h^{(T)}\partial_{\mu}h^{(L)\mu}\\ &+4(\partial_{\mu}h^{(L)\mu})^{2})\Big{]}.\end{split}$ (103)
@@
$\begin{split}-\frac{1}{2}m^{2}&(4\partial_{\mu}\partial_{\nu}l^{\parallel}\partial^{\mu}\partial^{\nu}l^{\parallel}-2(\Box l^{\parallel})^{2})=-m^{2}(\Box l^{\parallel})^{2}\\ &=-m^{2}l^{\parallel}(\Box)^{2}l^{\parallel},\end{split}$ (105)
@@
$\begin{split}-\frac{1}{2}m^{2}&\left(h_{\mu\nu}h^{\mu\nu}-\frac{1}{2}h^{2}\right)=-\frac{1}{2}m^{2}\Big{[}h_{\mu\nu}^{(T)}h^{(T)\mu\nu}-\frac{1}{2}h^{(T)2}\\ &-4h^{(L)\nu}\left(\partial^{\mu}h_{\mu\nu}^{(T)}-\frac{1}{2}\partial_{\nu}h^{(T)}\right)\\ &+2\partial_{\mu}l_{\nu}^{\perp}\partial^{\mu}l^{\perp\nu}-2m^{2}l^{\parallel}(\Box)^{2}l^{\parallel}\Big{]}\\ &=-\frac{1}{2}m^{2}\Big{[}h_{\mu\nu}^{(T)}h^{(T)\mu\nu}-\frac{1}{2}h^{(T)2}+2\partial_{\mu}l_{\nu}^{\perp}\partial^{\mu}l^{\perp\nu}\\ &+2l^{\parallel}(\Box)^{2}l^{\parallel}\Big{]},\end{split}$ (106)
@@
$\begin{split}S&=\int d^{4}x[\mathscr{L}_{FP}+\mathscr{L}_{1}],\ \mathscr{L}_{1}=-\frac{1}{4}m^{2}h^{2}\\ &=\int d^{4}x\Big{[}-\frac{1}{2}\partial_{\lambda}h_{\mu\nu}\partial^{\lambda}h^{\mu\nu}+\partial_{\mu}h_{\nu\lambda}\partial^{\nu}h^{\mu\lambda}-\partial_{\mu}h^{\mu\nu}\partial_{\nu}h\\ &+\frac{1}{2}\partial_{\lambda}h\partial^{\lambda}h-\frac{1}{2}m^{2}\left(h_{\mu\nu}h^{\mu\nu}-h^{2}\right)-\frac{1}{4}m^{2}h^{2}\Big{]}\,.\end{split}$ (107)
@@
$\begin{split}\mathscr{L}_{1}&=-\frac{1}{4}m^{2}\Big{[}h^{2}+4h\partial_{\mu}V^{\mu}+4(\partial_{\mu}V^{\mu})^{2}+4h\Box\phi\\ &+8\partial_{\mu}V^{\mu}\Box\phi+4(\Box\phi)^{2}\Big{]}.\end{split}$ (109)
@@
$\begin{split}\mathscr{L}_{1}&=-\frac{1}{4}m^{2}h^{2}-mh\partial_{\mu}V^{\mu}-(\partial_{\mu}V^{\mu})^{2}-h\Box\phi\\ &-2\frac{1}{m}\partial_{\mu}V^{\mu}\Box\phi-\frac{1}{m^{2}}(\Box\phi)^{2}.\end{split}$ (110)
@@
$\begin{split}\mathscr{L}_{1}&=-\frac{1}{4}m^{2}h^{{}^{\prime}2}-mh^{\prime}\partial_{\mu}V^{\mu}-(\partial_{\mu}V^{\mu})^{2}-h^{\prime}\Box\phi-2\frac{1}{m}\partial_{\mu}V^{\mu}\Box\phi\\ &-\frac{1}{m^{2}}(\Box\phi)^{2}-2m^{2}h^{\prime}\pi-4m^{2}\pi^{2}-4m\pi\partial_{\mu}V^{\mu}-4\pi\Box\phi.\end{split}$ (111)
@@
$\begin{split}\mathscr{L}_{1}&=-\frac{1}{4}m^{2}h^{{}^{\prime}2}-mh^{\prime}\partial_{\mu}V^{\mu}-(\partial_{\mu}V^{\mu})^{2}-h^{\prime}\Box\phi-2\frac{1}{m}\partial_{\mu}V^{\mu}\Box\phi\\ &-\frac{1}{m^{2}}(\Box\phi)^{2}-2m^{2}h^{\prime}\phi-4m^{2}\phi^{2}-4m\phi\partial_{\mu}V^{\mu}-4\phi\Box\phi.\end{split}$ (112)
@@
$\begin{split}S&=\int d^{4}x\Big{[}\mathscr{L}_{m=0}[h^{\prime}]-\frac{1}{2}m^{2}(h_{\mu\nu}^{\prime}h^{{}^{\prime}\mu\nu}-h^{{}^{\prime}2})-\frac{1}{2}F_{\mu\nu}F^{\mu\nu}\\ &+\left(-\phi\Box\phi+2m^{2}\phi^{2}\right)-2m(h_{\mu\nu}^{\prime}\partial^{\mu}V^{\nu}-h^{\prime}\partial_{\mu}V^{\mu})\\ &+(m^{2}h^{\prime}\phi+2m\phi\partial_{\mu}V^{\mu})+\kappa h_{\mu\nu}^{\prime}T^{\mu\nu}+\kappa\phi T\\ &-\frac{1}{4}m^{2}h^{{}^{\prime}2}-mh^{\prime}\partial_{\mu}V^{\mu}-(\partial_{\mu}V^{\mu})^{2}-h^{\prime}\Box\phi-2\frac{1}{m}\partial_{\mu}V^{\mu}\Box\phi\\ &-\frac{1}{m^{2}}(\Box\phi)^{2}\Big{]}\,.\end{split}$ (113)
@@
$\begin{split}S&+S_{GF1}+S_{GF2}=\int d^{4}x\Big{[}\mathscr{L}_{m=0}[h^{\prime}]-\frac{1}{2}m^{2}(h_{\mu\nu}^{\prime}h^{{}^{\prime}\mu\nu}-h^{{}^{\prime}2})\\ &-\frac{1}{2}F_{\mu\nu}F^{\mu\nu}+\left(-\phi\Box\phi+3m^{2}\phi^{2}\right)-\left(\partial^{\nu}h_{\mu\nu}^{\prime}-\frac{1}{2}\partial_{\mu}h^{\prime}\right)^{2}\\ &-m^{2}V_{\mu}V^{\mu}+\kappa h_{\mu\nu}^{\prime}T^{\mu\nu}+\phi T\Big{]}\\ &=\int d^{4}x\Big{[}\frac{1}{2}h_{\mu\nu}^{\prime}(\Box-m^{2})h^{{}^{\prime}\mu\nu}-\frac{1}{4}h^{\prime}(\Box-m^{2})h^{\prime}\\ &+V_{\mu}(\eta^{\mu\nu}\Box-\partial^{\mu}\partial^{\nu}-m^{2}\eta^{\mu\nu})V_{\nu}\\ &-3\phi(\Box-m^{2})\phi+\kappa h_{\mu\nu}^{\prime}T^{\mu\nu}+\kappa\phi T\Big{]}.\end{split}$ (115)
The relative sign of the first and third terms in the last line of the above equation does not change the equations of motion. Therefore, classically, such a change in sign is valid. Moreover, the above alterations of two signs do not change the theory even at the quantum level since all the correlation functions in quantum field theory that come from a single generating functional remain unchanged. The above action suggests no ghost DoF for $\alpha=1/2$ in four dimensions. However, for an arbitrary $\alpha$ in general $D$-dimensional space-time, the coefficients of each term may not be such that we can absorb the $(\Box\phi)^{2}$ in the gauge-fixing action.
Appendix D Degravitation
As mentioned in the main text, the mass of spin-2 field can be interpreted as a high-pass filter that sources must pass through before being detected by the spin-2 field [66, 64, 127]. This is analogous to the propagation of electromagnetic waves in plasma. In unmagnetized neutral plasma, the dispersion relation for electromagnetic waves is $\omega^{2}=k^{2}+\omega_{p}^{2}$, where $\omega_{p}$ is plasma frequency. Thus, waves with frequencies lower than $\omega_{p}$ cannot propagate in plasma. In the case of massive gravity, the short wavelength has little effect. As the cosmological constant is inherently a constant, it operates like a source with an exceedingly long wavelength. Thus, the cosmological constant acts as a screen that serves as a high-pass filter, leading to a considerably diminished observed effective value. This process is commonly referred to as degravitation [66, 64, 127].